OpenStudy (anonymous):

The total cost for 8 bracelets, including shipping was $54. The shipping charge was $6. Determine the cost for each bracelet. Show your work

OpenStudy (anonymous):

shipping charge=$6 cost of 8 bracle including shipping charge =$54 except shipping charge=54-6=$48 coast of each baraclet=48/8=$6 ans